OpenStudy (idealist10):

How to simplify e^(-1/2*cos(2x))?

ganeshie8 (ganeshie8):

it looks simplified to me ? @Kainui

OpenStudy (idealist10):

Then how to simplify (e^(-sin^2 x))*e^(-1/2*cos(2x))?

ganeshie8 (ganeshie8):

use below exponent property : \[\huge a^m*a^n = a^{m\color{red}{+}n}\]

OpenStudy (idealist10):

So e^(-sin^2 x-1/2*cos(2x))?

ganeshie8 (ganeshie8):

yes! \[\large e^{-\sin^2 x}*e^{-1/2*\cos(2x)} = e^{-\sin^2 x-1/2*\cos(2x)} \]

ganeshie8 (ganeshie8):

next, use the identity : \(\large \color{red}{\cos(2x) = 2\cos^2x - 1}\)

ganeshie8 (ganeshie8):

\[\large \begin{align} \\ e^{-\sin^2 x}*e^{-1/2*\cos(2x)} &= e^{-\sin^2 x-1/2*\cos(2x)} \\ &= e^{-\sin^2 x-1/2*(\color{Red}{2\cos^2x-1})} \\ \end{align} \]

ganeshie8 (ganeshie8):

\[\large \begin{align} \\ e^{-\sin^2 x}*e^{-1/2*\cos(2x)} &= e^{-\sin^2 x-1/2*\cos(2x)} \\ &= e^{-\sin^2 x-1/2*(\color{Red}{2\cos^2x-1})} \\ &= e^{-\sin^2 x-\color{Red}{\cos^2x+1/2}} \\ &= e^{-(\sin^2 x+\color{Red}{\cos^2x})\color{red}{+1/2}} \\ \end{align} \]

OpenStudy (idealist10):

So the final answer is e^-1/2?

ganeshie8 (ganeshie8):

yes ! if your prof hates negative fractions, you may give him : \(\large \dfrac{1}{\sqrt{e}}\)

OpenStudy (idealist10):

And how to integrate e^(-1/2) or 1/sqrt(e)?

ganeshie8 (ganeshie8):

\(e\) is just a number like \(1\),\(2\) or \(\pi\)

ganeshie8 (ganeshie8):

it is a constant.

ganeshie8 (ganeshie8):

\[\large \int k ~dx = kx+C\]

ganeshie8 (ganeshie8):

so, \[\large \int e^{-1/2}dx = e^{-1/2} \int dx = e^{-1/2} x + C \]

OpenStudy (idealist10):

So it's 1/sqrt(e)*x+C

ganeshie8 (ganeshie8):

yep !

OpenStudy (idealist10):

But what's (1/sqrt(e)*x+C)/(e^(-1/2*cos(2x)))?

ganeshie8 (ganeshie8):

whats the exact+complete question you're working on ?

OpenStudy (idealist10):

Find the general solution of y'+(2sin(x)cos(x))y=e^(-sin^2 x).

ganeshie8 (ganeshie8):

I see :) so you have got : \[\large \left(ye^{-1/2\cos(2x)}\right)' = e^{-\sin^2x}*e^{-1/2\cos(2x)}\]

ganeshie8 (ganeshie8):

after simplifying the right hand side, you get : \[\large \left(ye^{-1/2\cos(2x)}\right)' = e^{-1/2}\]

OpenStudy (idealist10):

maybe I also need to convert cos(2x)=2cos^2 x-1 again

ganeshie8 (ganeshie8):

then you integrate both sides : \[\large \int \left(ye^{-1/2\cos(2x)}\right)' dx = \int e^{-1/2}dx\]

ganeshie8 (ganeshie8):

and you arrive at : \[\large ye^{-1/2\cos(2x)} = e^{-1/2}x+C\]

ganeshie8 (ganeshie8):

right ?

OpenStudy (idealist10):

Yes, let me work it out.

ganeshie8 (ganeshie8):

there is a trick here to get the solution to a nice looking form

ganeshie8 (ganeshie8):

nvm, just simplify and see what u get

OpenStudy (idealist10):

I got e^(-cos^2 x+1/2)y=1/sqrt(e)*x+C

ganeshie8 (ganeshie8):

wolfram is giving this : http://prntscr.com/4bgh28

ganeshie8 (ganeshie8):

usually we leave the answer in implicit form unless somebody asks us to isolate y

ganeshie8 (ganeshie8):

\[\large ye^{-1/2\cos(2x)} = e^{-1/2}x+C\] you may leave the solution like this, no need to isolate y

OpenStudy (idealist10):

Thank you!